# ERPoT: Effective and Reliable Pose Tracking for Mobile Robots Based on Lightweight and Compact Polygon Maps

### v1

@@ Abstract
This paper presents an effective and reliable pose tracking solution termed ERPoT for mobile robots operating in large-scale outdoor environments, underpinned by an innovative prior polygon map. Especially, to overcome the challenge that arises as the map size grows with the expansion of the environment, the novel form of a prior map composed of multiple polygons is proposed. Benefiting from the use of polygons to concisely and accurately depict environmental occupancy, the prior polygon map achieves long-term reliable pose tracking while ensuring a compact form. More importantly, pose tracking is carried out under pure LiDAR mode, and the dense 3D point cloud is transformed into a sparse 2D scan through ground removal and obstacle selection. On this basis, a novel cost function for pose estimation through point-polygon matching is introduced, encompassing two distinct constraint forms: point-to-vertex and point-to-edge. In this study, our primary focus lies on two crucial aspects: lightweight and compact prior map construction, as well as effective and reliable robot pose tracking. Both aspects serve as the foundational pillars for future navigation across different mobile platforms equipped with different LiDAR sensors in different environments. Comparative experiments based on the publicly available datasets and our self-recorded datasets are conducted, and evaluation results show the superior performance of ERPoT on reliability, prior map size, pose estimation error, and runtime over the other five approaches. The corresponding code can be accessed at https://github.com/ghm0819/ERPoT, and the supplementary video is at https://youtu.be/cseml5FrW1Q.
@@ I Introduction
Motivated by the requirements mentioned above, we propose an effective and reliable pose tracking approach for mobile robots based on the novel prior map, called ERPoT. For mobile robots operating in terrestrial environments, the roll and pitch angles are typically stable and exhibit minimal variation [24]. Consequently, the proposed ERPoT, which operates on three degrees of freedom, is generally adequate to address the requirements of nearly all task scenarios effectively. In addition, the proposed pose tracking framework is a general solution that applies to various environments and is not limited to environments containing specific semantic elements. Specifically, LiDAR data serves as the sole input, from which a sparse 2D scan is derived through ground removal and obstacle selection. Subsequently, an offline map construction process is performed to acquire a lightweight and compact polygon map of the environment, as shown in Fig. 1. Multi-polygon-based representation (Fig. 1(c)) offers superior advantages over point cloud (Fig. 1(a)) and occupancy grid (Fig. 1(b)), particularly in applications demanding precision and efficient data utilization [25]. Finally, we can make use of the 2D scan and the prior polygon map to realize effective and reliable pose tracking for mobile robots, which takes the advantage of the newly proposed point-polygon matching. This research focuses on constructing lightweight and compact prior maps and establishing effective and reliable pose tracking. The main contributions include three aspects:
@@ II-B Prior map construction
The method and form of prior map construction are fundamental to the performance of prior map-based pose tracking for mobile robots. PoseMap proposed in [31] integrates the trajectory directly into the prior map, each robot pose is associated with a submap consisting of surfels, transcending the limitations of purely metric map representations and enhancing the overall efficiency. This topological structure of prior maps can improve pose tracking efficiency in large-scale environments. In order to fulfill the increasing computational requirements that come with the size of the prior map, Feng et al. [32] propose a localization system based on Block Maps to reduce the computational load. In [33], Xie et al. present a solution for robust indoor localization in environments like offices and corridors, utilizing 3D LiDAR point clouds within a compact, hierarchical, topometric semantic map.
@@ II-C Pose tracking
Accurate prior map-based pose tracking allows mobile robots to determine their exact location within an environment, which is essential for navigating and interacting with their surroundings effectively. Especially, for autonomous navigation, self-driving cars commonly rely on precise localization within a prior map to guide their way [37]. In [38], Wolcott et al. model the world as a mixture of several Gaussians and present a generic probabilistic-based localization system that can handle harsh weather conditions and poorly textured roadways. Based on the mesh-based prior map, the authors in [19] introduce an innovative observation model, seamlessly integrating it within a Monte Carlo localization framework to enhance accuracy and robustness. The Monte Carlo-based localization method has been widely used in large-scale outdoor tasks due to its advantages in dealing with uncertainty, nonlinear problems, and computational efficiency. For instance, Akai [20] merges the strengths of Monte Carlo localization with those of scan matching, culminating in the proposal of an efficient map-based localization solution that capitalizes on the best of both methodologies. Considering non-flat terrain characteristics of the operational environments, Rico et al. [21] present a mobile robot localization algorithm grounded in AMCL, utilizing elevation maps and probabilistic 3D occupancy maps to accurately determine the robot pose.
On the other hand, scan matching also plays a pivotal role in mobile robot pose tracking by aligning sequential point cloud data with the prior map to compute the robot pose. In [39], Koide et al. propose a real-time sensor localization system, integrating the NDT scan matching [11] with an angular velocity-based pose prediction using unscented Kalman filter (UKF). For long-term applications of mobile robots, the work in [40] introduces a resilient long-term LiDAR-based localization system, which incorporates a temporary mapping module to circumvent potential localization failures due to incorrect global matching. In contrast, in a data-driven manner, Hroob et al. [41] propose a novel stability scan filter for robust localization of field mobile robots through long-term stability landmarks in a continuously changing environment. In addition, the authors in [42] present a novel tightly-coupled graph localization approach that exploits prior topo-metric knowledge of the environment, and seamlessly integrates various modules into a graph optimization framework.
@@ III-A2 Online pose tracking
To reduce the size of 3D LiDAR points used for offline map construction and online pose tracking, the process of 2D scan generation is firstly carried out, which converts the dense 3D point cloud into a sparse 2D scan. This section describes the detailed process of 2D scan generation, as shown in Fig. 3.
@@
At first, the whole point cloud is projected into the fan-shaped grid map based on the concentric zone model, and for each region, the ground plane fitting is carried out to obtain the corresponding expression. Subsequently, all LiDAR points are segmented into two classes: ground points and obstacle points, and the obstacle point cloud is obtained through ground removal. Note that the detailed description of the ground segmentation process can refer to [44].
Considering that the workspace of mobile robots mainly lies on the ground, accurate and real-time three degrees of freedom pose estimation is enough for mobile robots to complete various complex tasks. Therefore, we will convert the dense 3D obstacle points into sparse but representative 2D obstacle points, which is enough for mobile robot pose tracking. On the other hand, different from the previous obstacle detection work [45], which focuses on the recall rate of obstacle points, while we are more concerned about making use of precise environmental information (obstacle points) to obtain accurate pose tracking results. Based on ground plane fitting for each region of the fan-shaped grid map, the 3D obstacle point in regions with unreliable ground fitting results can be removed for more accurate environmental information. In addition, the obstacle points higher than the mobile robot, defined as high-altitude points, are also unreliable and removed
Afterward, the valid 3D obstacle points are further converted into 2D obstacle points via ray-wise obstacle sorting and 2D obstacle generation. Specifically, let $\delta_{a}$ denote the horizontal angle resolution of the LiDAR sensor, and the corresponding number $N$ of 2D obstacle points is set as
$N=\lceil{2\pi/{\delta_{a}}}\rceil.$ (5)
Each 2D obstacle point is obtained according to the ray-wise obstacle sorting, and the number of rays is also $N$, while each ray contains an indefinite number of 3D obstacle points. For the $i$-th valid obstacle point $\mathbf{p}_{i}=[x_{i},y_{i},z_{i}]$, the range information $r_{i}$ can be calculated, as shown in Fig. 3(b). In addition, the horizontal angle $\theta_{i}$ and the corresponding ray index $j$ are calculated as follows
Based on the above, the $j$-th ray $\mathbf{R}_{j}$ is denoted by $\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{k}\}$ with $k$ being the number of obstacle points assigned into this ray. Then the final 2D obstacle point $\mathbf{p}_{i}^{2D}=[x_{i},y_{i}]$ is generated according to the nearest 3D obstacle point ($\mathbf{p}_{i}$ in Fig. 3(b)). If there is no valid obstacle point in $R_{i}$, then $\mathbf{p}_{i}^{2D}$ is set as $[r_{max}\cdot\cos(j\delta_{a}),r_{max}\cdot\sin(j\delta_{a})]$ with $r_{max}$ being the given maximum observation distance.
@@ III-C1 Grid map generation
Let $\mathbf{C}_{i}$ denote the 2D scan, which is represented by the set $\mathbf{C}_{i}=\{\mathbf{p}_{1}^{2D},\mathbf{p}_{2}^{2D},...,\mathbf{p}_{N}^{2D}\}$, corresponding to the sparse environment perception result of the $i$-th frame $\mathbf{l}_{i}$. For pose estimation, reliable static obstacles can provide more positive effects, however, there are several observation noises of the LiDAR sensor as well as dynamic obstacles in the environment (pedestrians, vehicles, and unstable vegetation). On the other hand, how to efficiently fuse multiple observations of the same scene, compactly realize the representation of the environment, and construct the prior map are also important issues.
@@
In order to address the above issues, as shown in Fig. 4, the occupancy grid map generation is introduced to fuse multiple observations and filter unstable obstacles, which is combined with the corresponding pose sequence ${{\cal S}_{t}}$. Therefore, equation (2) can be rewritten as
@@
where ${\cal C}_{t}=\{\mathbf{C}_{1},\mathbf{C}_{2},...,\mathbf{C}_{t}\}$ denotes the 2D scan sequence. In Fig. 4, the map origin is marked by a black filled circle, denoted by the offset information $(x_{c},y_{c})$ and finally determined only after the construction process of the grid map, and the resolution is denoted by $r$, which is the size of each grid.
Furthermore, the construction problem of grid map ${\cal M}_{g}$ is converted into the state estimation problem of each grid, which is denoted as
@@
where $g_{i}$ represents the $i$-th grid, and the number of grids is denoted by $N_{g}$. As shown in Fig. 4, we make use of the ray trace algorithm to update the grid map by connecting the scan point and the sensor origin. Specifically, the number of hits $N_{hit}$ and the number of passes $N_{pass}$ of every grid are separately counted, and the occupancy probability of the $i$-th grid $g_{i}$ is calculated as
@@
In the generation process, the size and the state of the grid map gradually update as new areas are observed. Finally, the occupancy probabilities of all grids are calculated, and each grid is determined whether the current state is occupied based on the given threshold.
@@ III-C2 Polygon map construction
After the above generation process of the grid map, the binary image representing the static obstacles in the environment can be obtained. On this basis, the contour extraction process of the grid map is carried out to achieve a lightweight and compact vectorized representation, as shown in Fig. 5.
@@
(2) Afterward, the process of contour extraction via the newly obtained binary map is obtained, and the results are shown in Fig. 5(c), which are represented by the red and directional contours.
(3) More importantly, the red contour does not match the actual obstacle information (Fig. 5(c)), and the final accurate contours are obtained via the process of contour translation, as shown in Fig. 5(d).
Furthermore, we take the points denoted by $\mathbf{p}_{i^{\prime}}^{c}$ and $\mathbf{p}_{i}^{c}$ as an example and describe how to obtain the coordinate information of the contour vertex. Specifically, the image coordinate for $\mathbf{p}_{i^{\prime}}^{c}$ is denoted by $(\text{row}_{i},\text{col}_{i})$, representing the indexes of the row and the column, respectively. Therefore, in the grid map coordinate system, $\mathbf{p}_{i^{\prime}}^{c}=\left[x_{i^{\prime}}^{c},y_{i^{\prime}}^{c}\right]$ is calculated as follows
@@
In particular, as depicted in Fig. 6(a), the example of polygon segmentation involves a contour with dozens of vertices arranged in counter-clockwise order. In the initial segmentation step, the vertices denoted as $\mathbf{p}_{10}^{w}$ and $\mathbf{p}_{34}^{w}$ serve as endpoints of the segmentation line, effectively dividing the contour into two roughly equal parts. This result is illustrated in Fig. 6(b). Ultimately, following two iterations of polygon segmentation, the original contour is divided into four distinct polygons, as illustrated in Fig. 6(d). After polygon segmentation and simplification, we make use of the polygon map ${\cal M}_{p}=\{\mathbf{P}_{1},\mathbf{P}_{2}\,...,\mathbf{P}_{m}\}$ to represent the environmental occupancy, and the $i$-th $\mathbf{P}_{i}$ is also denoted by the vertex set $\{\mathbf{p}_{1}^{w},\mathbf{p}_{2}^{w},,...,\mathbf{p}_{N_{i}}^{w}\}$ with $N_{i}$ being the number of vertices.
@@ III-D Online pose tracking
Once the polygon map ${\cal M}_{p}$ of the environment has been obtained, it can serve as prior information for effective and reliable pose tracking. Specifically, the newly proposed pose tracking approach encompasses the following detailed processes: feature extraction, point-polygon matching, and real-time pose tracking.
@@ III-D1 Feature extraction
Afterward, according to the given smoothness threshold, the corner feature set $\mathbb{F}_{c}$ and edge feature set $\mathbb{F}_{e}$ can be obtained. Specifically, we exclusively employ the derived single-line scan for feature extraction, while corner features and edge features are identified as scan points exhibiting relatively lower and higher smoothness values, respectively, as shown in Fig. 7. Note that the obtained nearest obstacle points (Fig. 7(b)) retain their x-coordinates and y-coordinates, thereby compressing them into a 2D scan (Fig. 7(c)) for feature extraction, which is different from the conventional methodologies employed in 3D LiDAR odometry.

### v3

# ERPoT: Effective and Reliable Pose Tracking for Mobile Robots Using Lightweight Polygon Maps

@@ Abstract
This paper presents an effective and reliable pose tracking solution, termed ERPoT, for mobile robots operating in large-scale outdoor and challenging indoor environments, underpinned by an innovative prior polygon map. Especially, to overcome the challenge that arises as the map size grows with the expansion of the environment, the novel form of a prior map composed of multiple polygons is proposed. Benefiting from the use of polygons to concisely and accurately depict environmental occupancy, the prior polygon map achieves long-term reliable pose tracking while ensuring a compact form. More importantly, pose tracking is carried out under pure LiDAR mode, and the dense 3D point cloud is transformed into a sparse 2D scan through ground removal and obstacle selection. On this basis, a novel cost function for pose estimation through point-polygon matching is introduced, encompassing two distinct constraint forms: point-to-vertex and point-to-edge. In this study, our primary focus lies on two crucial aspects: lightweight and compact prior map construction, as well as effective and reliable robot pose tracking. Both aspects serve as the foundational pillars for future navigation across diverse mobile platforms equipped with different LiDAR sensors in varied environments. Comparative experiments based on the publicly available datasets and our self-recorded datasets are conducted, and evaluation results show the superior performance of ERPoT on reliability, prior map size, pose estimation error, and runtime over the other six approaches. The corresponding code can be accessed at https://github.com/ghm0819/ERPoT, and the supplementary video is at https://youtu.be/6XdcXyUrLKw.
@@ I Introduction
Motivated by the requirements mentioned above, we propose an effective and reliable pose tracking approach for mobile robots based on the novel prior map, called ERPoT. For mobile robots operating in terrestrial environments, the roll and pitch angles are typically stable and exhibit minimal variation [24]. Consequently, the proposed ERPoT, which operates on three degrees of freedom, is generally adequate to effectively address the requirements of nearly all task scenarios. In addition, the proposed pose tracking framework is a general solution that applies to various environments and is not limited to environments containing specific semantic elements. Specifically, LiDAR data serves as the sole input, from which a sparse 2D scan is derived through ground removal and obstacle selection. Subsequently, an offline map construction process is performed to acquire a lightweight and compact polygon map of the environment, as shown in Fig. 1. Multi-polygon-based representation (Fig. 1(c)) offers superior advantages over point cloud (Fig. 1(a)) and occupancy grid (Fig. 1(b)), particularly in applications demanding precision and efficient data utilization [25]. Finally, we can make use of the 2D scan and the prior polygon map to realize effective and reliable pose tracking for mobile robots, which takes the advantage of the newly proposed point-polygon matching. This research focuses on constructing lightweight and compact prior maps and establishing effective and reliable pose tracking. The main contributions include three aspects:
@@ II-B Prior map construction
The method and form of prior map construction are fundamental to the performance of prior map-based pose tracking for mobile robots. PoseMap proposed in [31] integrates the trajectory directly into the prior map, each robot pose is associated with a submap consisting of surfels, transcending the limitations of purely metric map representations and enhancing the overall efficiency. This topological structure of prior maps can improve pose tracking efficiency in large-scale environments. To fulfill the increasing computational requirements that come with the size of the prior map, Feng et al. [32] propose a localization system based on Block Maps to reduce the computational load. In [33], Xie et al. present a solution for robust indoor localization in environments like offices and corridors, utilizing 3D LiDAR point clouds within a compact, hierarchical, topometric semantic map.
@@ II-C Pose tracking
Accurate prior map-based pose tracking allows mobile robots to determine their exact location within an environment, which is essential for navigating and interacting with their surroundings effectively. Especially, for autonomous navigation, self-driving cars commonly rely on precise localization within a prior map to guide their way [37]. In [38], Wolcott et al. model the world as a mixture of several Gaussians and present a generic probabilistic-based localization system that can handle harsh weather conditions and poorly textured roadways. Based on the mesh-based prior map, the authors in [19] introduce an innovative observation model, seamlessly integrating it within a Monte Carlo localization framework to enhance accuracy and robustness. The Monte Carlo-based localization method has been widely used in large-scale outdoor tasks due to its advantages in dealing with uncertainty, nonlinear problems, and computational efficiency. For instance, Akai [20] merges the strengths of Monte Carlo localization with those of scan matching, culminating in the proposal of an efficient map-based localization solution that capitalizes on the best of both methodologies. Considering non-flat terrain characteristics of the operational environments, Rico et al. [21] present a mobile robot localization algorithm grounded in AMCL, utilizing elevation maps and probabilistic 3D occupancy maps to determine the robot pose accurately.
On the other hand, scan matching also plays a pivotal role in mobile robot pose tracking by aligning sequential point cloud data with the prior map to compute the robot pose. In [39], Koide et al. propose a real-time sensor localization system, integrating the NDT scan matching [11] with an angular velocity-based pose prediction using an unscented Kalman filter (UKF). For long-term applications of mobile robots, the work in [40] introduces a resilient long-term LiDAR-based localization system, which incorporates a temporary mapping module to circumvent potential localization failures due to incorrect global matching. In contrast, in a data-driven manner, Hroob et al. [41] propose a novel stability scan filter for robust localization of field mobile robots through long-term stability landmarks in a continuously changing environment. In addition, the authors in [42] present a novel tightly-coupled graph localization approach that exploits prior topo-metric knowledge of the environment and seamlessly integrates various modules into a graph optimization framework.
@@ III-A2 Online pose tracking
Considering that the workspace of mobile robots mainly lies on the ground, accurate and real-time three degrees of freedom pose estimation is enough to complete various complex tasks. Therefore, we convert the dense 3D obstacle points into sparse but representative 2D obstacle points, which is enough for mobile robot pose tracking. This section describes the detailed process of 2D scan generation, as shown in Fig. 3.
At first, the point cloud is projected into a fan-shaped grid map using the concentric zone model. Ground plane fitting [44] is subsequently performed for each region to separate ground and obstacle points. Unlike previous work [45] that emphasized obstacle point recall, we prioritize accurate pose tracking using precise environmental information. Unreliable obstacle points in regions with poor ground fitting and high-altitude points are removed to enhance accuracy.
Afterward, the valid 3D obstacle points are transformed into 2D obstacle points through ray-wise sorting and 2D generation. Specifically, the number $N$ of 2D obstacle points is determined by the LiDAR’s horizontal angle resolution $\delta_{a}$,
$N=\lceil{2\pi/{\delta_{a}}}\rceil.$ (5)
Accordingly, 2D obstacle points are derived from ray-wise sorting, with $N$ rays, each having multiple 3D points. For the $i$-th valid point $\mathbf{p}_{i}=[x_{i},y_{i},z_{i}]$, the range $r_{i}$ is calculated as shown in Fig. 3(b), and the horizontal angle $\theta_{i}$ and ray index $j$ are determined as follows
Based on the above, the ray $\mathbf{R}_{j}$ is denoted by $\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{k}\}$ with $k$ being the number of points. After nearest obstacle selection, the final 2D point $\mathbf{p}_{i}^{2D}=[x_{i},y_{i}]$ is based on the nearest 3D point ($\mathbf{p}_{i}$ in Fig. 3(b)).
@@ III-C1 Grid map generation
Let $\mathbf{C}_{i}$ denote the 2D scan, represented by $\mathbf{C}_{i}=\{\mathbf{p}_{1}^{2D},\mathbf{p}_{2}^{2D},...,\mathbf{p}_{N}^{2D}\}$, corresponding to the sparse perception. While reliable static obstacles aid pose estimation, LiDAR noise and dynamic obstacles pose challenges. Additionally, efficiently fusing multiple observations, compactly representing the environment, and constructing the prior map are crucial tasks.
@@
Accordingly, the occupancy grid map generation is introduced in Fig. 4, which is combined with the corresponding pose sequence ${{\cal S}_{t}}$. Therefore, equation (2) can be rewritten as
@@
where ${\cal C}_{t}=\{\mathbf{C}_{1},\mathbf{C}_{2},...,\mathbf{C}_{t}\}$ denotes the 2D scan sequence. The construction problem of grid map ${\cal M}_{g}$ is converted into the state estimation problem of each grid as follows
@@
where $g_{i}$ is the $i$-th grid, and $N_{g}$ is the total number of grids. As shown in Fig. 4, the ray trace algorithm updates the grid map by connecting scan points to the sensor origin. Specifically, each grid’s hits $N_{hit}$ and passes $N_{pass}$ are counted, and the occupancy probability of grid $g_{i}$ is calculated as
@@
In the generation process, the size and the state of the grid map gradually update as new areas are observed. Finally, the occupancy probabilities of all grids are calculated, and each grid determines whether the current state is occupied based on the given threshold.
@@ III-C2 Polygon map construction
After the above generation process of the grid map, the binary image representing the static obstacles in the environment can be obtained. On this basis, the contour extraction process of the grid map is carried out to achieve a lightweight and compact representation.
@@
(2) Subsequently, the contour extraction process using the newly obtained binary map is performed, with the results depicted in Fig. 5(c) as red and directional contours.
(3) Moreover, the red contour in Fig. 5(c) does not match the actual obstacle information. The accurate final contours are obtained through contour translation, as shown in Fig. 5(d).
Furthermore, using points $\mathbf{p}_{i^{\prime}}^{c}$ and $\mathbf{p}_{i}^{c}$ as an example, we explain how to obtain the contour vertex coordinates. Specifically, the image coordinates of $\mathbf{p}_{i^{\prime}}^{c}$ are $(\text{row}_{i},\text{col}_{i})$, representing the row and column indexes. In the grid map coordinate system, $\mathbf{p}_{i^{\prime}}^{c}=\left[x_{i^{\prime}}^{c},y_{i^{\prime}}^{c}\right]$ is calculated as follows
@@
In particular, as depicted in Fig. 6(a), the example of polygon segmentation involves a contour with dozens of vertices arranged in counter-clockwise order. In the initial segmentation step, the vertices denoted as $\mathbf{p}_{10}^{w}$ and $\mathbf{p}_{34}^{w}$ serve as endpoints of the segmentation line, effectively dividing the contour into two roughly equal parts. This result is illustrated in Fig. 6(b). Ultimately, following two iterations of polygon segmentation, the original contour is divided into four distinct polygons, as illustrated in Fig. 6(d). After polygon segmentation and simplification, we make use of the polygon map ${\cal M}_{p}=\{\mathbf{P}_{1},\mathbf{P}_{2}\,...,\mathbf{P}_{m}\}$ to represent the environmental occupancy, and the $i$-th $\mathbf{P}_{i}$ is also denoted by the vertex set $\{\mathbf{p}_{1}^{w},\mathbf{p}_{2}^{w},,...,\mathbf{p}_{N_{i}}^{w}\}$ with $N_{i}$ being the number of vertices. Note that throughout the entire process of polygon segmentation and subsequent pose tracking, the counter-clockwise characteristic of the polygon vertices is preserved, as shown in Fig. 6.
III-C3 Prior map representation
Given the polygon map ${\cal M}_{p}$, before commencing with pose tracking, the process of prior map representation is carried out, and this step is crucial for establishing a solid foundation for accurate tracking. Specifically, for the $i$-th polygon $\mathbf{P}_{i}=\{\mathbf{p}_{1}^{w},\mathbf{p}_{2}^{w},...,\mathbf{p}_{N_{i}}^{w}\}$, the corresponding centroid point $\mathbf{p}_{i}^{e}=[x_{i}^{e},y_{i}^{e}]$ is calculated as
$x_{i}^{e}=\frac{m_{10}}{m_{00}},\ y_{i}^{e}=\frac{m_{01}}{m_{00}},$ (15)
@@ III-D Online pose tracking
Once the prior map of the environment has been obtained, it can be utilized as prior information to facilitate effective and reliable pose tracking. Specifically, the newly proposed pose tracking approach encompasses the following detailed processes: feature extraction, point-polygon matching, and real-time pose tracking.
@@ III-D1 Feature extraction
Afterward, according to the given smoothness threshold ($\sigma_{c}$ and $\sigma_{e}$), the corner feature set $\mathbb{F}_{c}$ and edge feature set $\mathbb{F}_{e}$ can be obtained. Specifically, we exclusively employ the derived single-line scan for feature extraction, while corner features and edge features are identified as scan points exhibiting relatively lower and higher smoothness values, respectively, as shown in Fig. 7. Note that the obtained nearest obstacle points (Fig. 7(b)) retain their x-coordinates and y-coordinates, thereby compressing them into a 2D scan (Fig. 7(c)) for feature extraction, which is different from the conventional methodologies employed in 3D LiDAR odometry.